# LoQT: Low Rank Adapters for Quantized Training

### v1

@@ B.2 Fine-tuning
We test learning rates in the range of $1\times 10^{-5}$ to $5\times 10^{-4}$. For LoQT qLoFT, we employed normal float NF4 quantization and performed 5 iterations of optimizing the error of quantization. We used a batch size of 32 and a maximum sequence length of 256. Table 5 summarizes the detailed hyperparameters for tasks in GLUE using the DeBERTaV3-base model. We use a fixed projection gap of 2400 for all runs.
@@ Appendix C Rank Ablation
Figure 5 shows the validation perplexity versus training steps for various ranks using LoQT-nq and LoQT on a $130$ million parameter model over 20,000 iterations. All models employ an exponentially increasing update frequency starting at $100$, with a factor of $1.2^{T_{i}}$. The results demonstrate that both the quantized (LoQT) and non-quantized (LoQT-nq) models follow a similar trajectory for ranks ranging from $64$ to $512$. However, for the smaller rank of $64$, there is a slight divergence between LoQT-nq and LoQT, indicating a limit to how low the rank can be while maintaining accuracy with quantization. This plot highlights the tradeoff between rank and perplexity, suggesting that while our method supports low-rank training, there is a minimum rank threshold needed to achieve results comparable to regular pre-training.
@@ C.1 Memory Measurements
Figure 6 demonstrates that LoQT requires less memory than GaLore and Adam, even without using per-layer gradients [25] or Adam 8-bit [5]. The gap between LoQT and the baselines increases with larger model sizes. The configurations and ranks for each model are shown in Table 4. With LoQT and Adam 8-bit, it is possible to pre-train a 13B model with rank 1024 on a GPU with 24GB of VRAM. This enables training with LoQT on consumer GPUs, such as the NVIDIA 4090, using a small per-GPU token batch size 256. Figure 1 in the main text provides a detailed breakdown of each memory component for the 13B model. Maximum memory allocated is measured using nvitop (https://github.com/XuehaiPan/nvitop).
@@
Figure 6: Memory usage for LoQT vs baselines for different model sizes. LW means per-layer gradient updates as per [25], and A8bit mean with Adam 8-bit. We evaluate using a token batch size of 256.

### v4

@@ B.2 Fine-tuning
We test learning rates in the range of $1\times 10^{-5}$ to $5\times 10^{-4}$. For LoQT and LoftQ, we employed normal float NF4 quantization and performed five iterations of optimizing the error of quantization. We used a batch size of 32 and a maximum sequence length of 256. Tab. 7 summarizes the detailed hyperparameters for tasks in GLUE using the DeBERTaV3-base model. We use a fixed projection gap of 2400 for all runs. Each of the parameter-efficient training methods is applied to all linear layers of the network, including attention projection and feed-forward layers, while the embedding layer is not trained.
@@ Appendix C Rank Ablation
Fig. 5 shows the validation perplexity versus training steps for various ranks using LoQT-nq and LoQT on a $130$ million parameter model over 20,000 iterations. All models employ an exponentially increasing update frequency starting at $100$, with a factor of $1.2^{T_{i}}$. The results demonstrate that both the quantized (LoQT) and non-quantized (LoQT-nq) models follow a similar trajectory for ranks ranging from $64$ to $512$. However, for the smaller rank of $64$, there is a slight divergence between LoQT-nq and LoQT, indicating a limit to how low the rank can be while maintaining accuracy with quantization. This plot highlights the tradeoff between rank and perplexity, suggesting that while our method supports low-rank training, there is a minimum rank threshold needed to achieve results comparable to regular pretraining.
@@ C.1 Memory Measurements
Fig. 6 demonstrates that LoQT requires less memory than GaLore and Adam, even without using per-layer gradients [13] or Adam 8-bit [20]. The gap between LoQT and the baselines increases with larger model sizes. The configurations and ranks for each model are shown in Tab. 6. With LoQT and Adam 8-bit, it is possible to pretrain a 13B model with rank 1024 on a GPU with 24GB of VRAM. This enables training with LoQT on consumer GPUs, such as the NVIDIA 4090, using a small per-GPU token batch size of 256. Fig. 1 in the main text provides a detailed breakdown of each memory component for the 13B model. Maximum memory allocated is measured using nvitop (https://github.com/XuehaiPan/nvitop).
@@
Figure 6: Memory usage for LoQT vs baselines for different model sizes. LW means per-layer gradient updates as per [13], and A8bit means with Adam 8-bit. We evaluate using a token batch size of 256.
Finetuning without merging low-rank factors
Tab. 9 shows how LoQT and LoQT-nq perform when not using the merging of factors while training. We see that the results are slightly worse than those where merging is performed; however, the results are still better than LoftQ, LoRA, and Galore, showing that in the finetuning case, where we already have a pretrained model, we can omit to merge low-rank factors into W, and still get good results. This would allow W to be kept fixed and quantized while training only the adapters, which enables using LoQT like LoRA, where multiple adapters can be trained for the same set of frozen weights.
Task adaptation for GSM8K on Llama 7B and 13B
The GSM8K dataset [12] "is a dataset of 8.5K high-quality linguistically diverse grade school math word problems created by human problem writers. The dataset is segmented into 7.5K training problems and 1K test problems". It has been used extensively to benchmark LLMs and is a good candidate for evaluating LoQT. We experimented with 7B and 13B models, performing a hyperparameter search to find the optimal learning rate for each method. Using the best learning rate, we trained each model over three seeds for three epochs with a sequence length of 512, applying 4-bit quantization for fine-tuning Llama-2 models on the GSM8K training set. We report the average test set accuracy and standard error in Tab. 4. LoQT achieves an accuracy of 42.6 for Llama-7B and 52.9 for Llama-2 13B. Both results are average obtained over three seeds without merging and with rank 64. Tab. 8 lists the hyper-parameters. We evaluate the fine-tuned and quantized model on the validation set and report the best perplexity across learning rates in Tab. 8. For each of the methods, only the attention projection matrices are trained.
Table 8: Hyper-parameters used for the GSM8K task.
Hyper-parameter Value
Optimizer AdamW
Weight decay 0.1
LR $\{0.1,0.5,0.7,1,3,4\}\times 10^{-4}$
LR scheduler cosine
Warmup ratio 3%
Epochs 3
Batch size 16 (7B), 8 (13B)
Max sequence length 512
Continued pretraining of Llama 7B
We run our pretraining experiments for models up to 1B parameters. To demonstrate the feasibility of LoQT on a larger model size in a practical application, we do continued pretraining for language adaptation. To this end, we start with the meta-llama
Llama-2-7b-hf
from Hugging face and run 1500 updates (1 epoch) on the model using a batch size of 512. We compare LoQT with NF4 quantization, LoQT-nq (no quantization), regular training, and GaLore. We use rank 1024 for all models where applicable, adam8bit optimizer, and 150 warmup steps. For LoQT and GaLore we use a learning rate of 0.0002 with a scale factor of 0.25 and for the others a learning rate of 0.00005. The dataset we train on is a curated subset of a public Icelandic text dataset [10] with  770k documents and a corresponding evaluation set. We released the data splits at (https://huggingface.co/datasets/vesteinn/loqt_icelandic). We chose Icelandic since the model has a limited performance on the language yet it was included to some degree in the pretraining data, enabling a clear improvement trajectory. The results comparing Galore, Regular training (Full), and LoQT are shown in Tab. 4. LoQT and LoQT-nq perform the best at 3.61 and 3.63 in perplexity, similar to full training at 3.79, while GaLore gets 3.96 and the original model 4.90.
Appendix D Memory-Saving Ablations
To evaluate the differences between memory savings with layer-wise gradient computations and 8-bit Adam, we conduct an ablation experiment using a 130M parameter model. We compare four settings: regular training, 8-bit Adam, layer-wise gradient updates, and a combination of 8-bit Adam with layer-wise updates, tested on Galore, LoQT, and regular FP16 training. Our results, illustrated in Fig. 9 show that adding these memory-saving components introduces a small decrease in performance. Importantly, LoQT experiences a proportionally smaller decrease in performance compared to Galore and full training when combining both 8-bit and layer-wise updates. These results demonstrate that while memory savings come with some trade-offs, LoQT maintains good performance. In addition, due to the lower memory requirements of LoQT, we enable training of larger models without using layer-wise gradient computations and 8-bit Adam.
Memory savings with varying sequence lengths
With larger contexts, the overall memory consumption is increasingly influenced by activations. Following prior work [8, 48], our experimentation has focused on the setting of shorter context lengths (256 tokens). But as demonstrated in Fig. 8, the benefit of LoQT does transfer to longer context lengths, enabling training of Llama 7B on consumer hardware with a context length of 2048 and 4096 on a 40GB A100 without activation checkpointing.
Appendix E Generalization to other architectures and models
LoQT should work with any type of deep neural network using linear layers, such as vision transformers or state space models. To narrow the scope of our work and provide a more detailed analysis, however, we choose to focus on a well-studied auto-regressive language model and a bi-directional masked language model that has been commonly used as a basis in much of the related work. We hope to see LoQT being used for other model architectures.
Table 9: Results with LoQT, LoQT-nq, and GaLore of DeBERTaV3-base models on the GLUE development set. We report mean and standard error over three seeds. The best results on each dataset are shown in bold. "No-Merge" means we do not update the pretrained matrix $W$ during training.
Rank Method MNLI QNLI RTE SST MRPC CoLA QQP STSB Average
Acc Acc Acc Acc f1 Matt f1 PCorr
32 LoQT-nq - no Merge 90.0$\pm$0.10 94.1$\pm$0.01 84.5$\pm$0.01 95.6$\pm$0.03 93.8$\pm$0.01 72.0$\pm$0.01 89.8$\pm$0.01 91.6$\pm$0.01 89.0
32 LoQT- no Merge 90.0$\pm$0.12 94.1$\pm$0.01 86.1$\pm$0.15 95.7$\pm$0.02 94.2$\pm$0.01 71.4$\pm$0.20 89.6$\pm$0.01 90.8$\pm$0.01 88.9
32 LoftQ 90.4$\pm$0.09 93.2$\pm$0.02 83.8$\pm$0.63 95.6 $\pm$0.07 93.2$\pm$0.14 71.1$\pm$0.28 89.6$\pm$0.12 91.0$\pm$0.09 88.4
Figure 7: Naive Quantization of W and P vs including Error Compensation(EC) and Exp. increasing intervals(EI).
Figure 8: Memory usage for common context lengths (256 to 4096) for the LLaMA 7B model measured using torch.cuda.max_memory_allocated. We include lines representing 16GB, 24GB, and 40GB VRAM limits to indicate which configurations fit within the VRAM capacities of standard NVIDIA GPUs.
Figure 9: Validation perplexity with various optimization techniques: Adam 8bit, per-layer updates, and their combinations, compared to baseline training without these optimizations. LW means per-layer gradient updates as per [13], and Adam8bit means with Adam 8-bit.